# UniGenX: Unified Generation of Sequence and Structure with Autoregressive Diffusion

### v1

@@ Abstract
Unified generation of sequence and structure for scientific data (e.g., materials, molecules, proteins) is a critical task. Existing approaches primarily rely on either autoregressive sequence models or diffusion models, each offering distinct advantages and facing notable limitations. Autoregressive models, such as GPT, Llama, and Phi-4, have demonstrated remarkable success in natural language generation and have been extended to multimodal tasks (e.g., image, video, and audio) using advanced encoders like VQ-VAE to represent complex modalities as discrete sequences. However, their direct application to scientific domains is challenging due to the high precision requirements and the diverse nature of scientific data. On the other hand, diffusion models excel at generating high-dimensional scientific data, such as protein, molecule, and material structures, with remarkable accuracy. Yet, their inability to effectively model sequences limits their potential as general-purpose multimodal foundation models. To address these challenges, we propose UniGenX, a unified framework that combines autoregressive next-token prediction with conditional diffusion models. This integration leverages the strengths of autoregressive models to ease the training of conditional diffusion models, while diffusion-based generative heads enhance the precision of autoregressive predictions. We validate the effectiveness of UniGenX on material and small molecule generation tasks, achieving a significant leap in state-of-the-art performance for material crystal structure prediction and establishing new state-of-the-art results for small molecule structure prediction, de novo design, and conditional generation. Notably, UniGenX demonstrates significant improvements, especially in handling long sequences for complex structures, showcasing its efficacy as a versatile tool for scientific data generation.
Keywords— Multimodality, Diffusion, Autoregressive, Science generation
@@ 1 Introduction
The ability to accurately model and generate both sequence and structure information within scientific data is a linchpin for advancing artificial intelligence in scientific discovery. This challenge, though critical, presents a significant hurdle, as the intricate interplay between sequence and structure often dictates the fundamental properties and functionalities of materials, molecules, and proteins. Structural information, therefore, serves as the bedrock for comprehending their behavior and designing transformative applications. For instance, the precise arrangement of atoms within a crystal lattice directly influences a material’s mechanical, thermal, and electronic characteristics [1], while the nuanced three-dimensional conformation of a protein governs its biological function through specific molecular interactions [2]. Similarly, the spatial configuration of small molecules exerts a profound influence on their chemical reactivity, pharmacological efficacy, and potential toxicity [3].
In these scientific domains, where precision is paramount, even subtle deviations in atomic coordinates or bond angles can precipitate substantial alterations in properties, rendering a material unsuitable for its intended application or disrupting a molecule’s interaction with critical biological targets [4]. For example, minor inaccuracies in molecular conformations can lead to erroneous predictions of binding affinities in drug discovery pipelines [5] or impede the precise replication of desired material properties in materials science laboratories [6]. Furthermore, the impact of structural inaccuracies is compounded when modeling larger, more complex systems, such as intricate protein complexes, extended periodic materials, or extensive molecular networks, where the complexity of the system amplifies the criticality of precision [7].
The inherent high-dimensional nature of scientific data further exacerbates the complexity of this modeling task. Unlike the symbolic or approximate representations often employed in natural language processing or image analysis, scientific data frequently relies on precise numerical representations, including three-dimensional atomic coordinates, lattice vectors, and energetic properties [8]. These representations are not only exceptionally sensitive to minor perturbations but also exhibit intricate, multiscale dependencies, spanning from atomic-level interactions to macroscopic properties. Consequently, the successful modeling of such data necessitates a profound understanding of both sequence and structure, coupled with the capacity to generate outputs that are both physically plausible and scientifically meaningful [9].
Adding to the complexity is the inherent diversity of scientific data formats across various domains. Materials are typically represented using chemical formulas and periodic lattice structures [10], molecules are characterized by SMILES strings and three-dimensional conformations [11], and proteins are described by amino acid sequences and spatial folding patterns [12]. Each of these representational paradigms encodes unique, complex relationships between sequence and structure, demanding specialized approaches to accurately capture their intricacies. The absence of a unified modeling framework capable of seamlessly handling this diversity significantly impedes the development of general-purpose artificial intelligence systems for scientific discovery.
Consequently, achieving the dual objectives of high precision and broad generalization in the modeling of sequence and structure remains a substantial bottleneck in the field of AI for science. Overcoming this challenge is crucial for advancing applications in diverse domains such as materials design, drug discovery, and protein engineering, where the ability to accurately model and predict the intricate interplay between sequence and structure can unlock transformative innovations [4, 13].
Autoregressive language models (LLMs), exemplified by models such as GPT [14], Llama [15], phi-4 [16], Gemini [17], and DeepSeek [18], have revolutionized the field of language generation and have been successfully extended to multimodal tasks, including image, video, and audio generation, through the utilization of vector quantized embeddings. These models benefit significantly from scaling laws, where increasing model size and data volume leads to improved performance. Their ability to process diverse data as token sequences, through direct tokenization for text or vector quantized embeddings for other modalities, confers both flexibility and generalizability. Furthermore, their capacity to handle long-context inputs enables the generation of complex outputs for both conditional and unconditional tasks. However, these models encounter significant challenges when applied to scientific domains, particularly in the generation of numerically accurate outputs, which is indispensable for tasks such as three-dimensional structure generation.
Conversely, diffusion models [19, 20, 21, 22] have demonstrated exceptional prowess in high-dimensional numerical generation tasks, achieving remarkable success in generating high-resolution images and videos [23, 24, 25, 26, 27]. These models typically encode high-dimensional data, such as images with dimensions of $512\times 512\times 3$, into lower-dimensional latent representations, such as $128\times 128\times h$ where $h$ is the hidden dimension, using encoders like variational autoencoders (VAEs). These representations are then patch-wisely tokenized, for example, $32\times 32$ with 1024 tokens, and processed by transformer-based models [28] to predict noise or score values. These tokens are subsequently treated as a concatenated high-dimensional vector ($1024\times h$), upon which the diffusion process is conditioned. While this approach enhances efficiency compared to directly diffusing in the original image space, training diffusion models remains computationally intensive and prone to defects [29], especially when not fully trained. Moreover, diffusion models generally struggle with sequence modeling tasks, limiting their applicability in scenarios requiring flexible multimodal capabilities.
The scientific domain, with its intricate numerical representations and diverse data formats, presents unique challenges for generative models. Even minute deviations in atomic coordinates can drastically alter molecular properties or render generated structures meaningless. While LLMs have shown promise in specific scientific applications [30, 31, 32, 33, 34, 35], they still struggle with unified numerical data modeling. Conversely, diffusion models are successful in generating structures [36, 37, 38, 39, 40, 41], but lack flexible multimodal capabilities.
To address these challenges, we introduce UniGenX, a novel framework integrating autoregressive next-token prediction with conditional diffusion models. This leverages autoregressive flexibility and diffusion precision, overcoming numerical accuracy limitations while maintaining multimodal capability.
Our approach offers several key advantages:
  • Improved numerical accuracy: The conditional diffusion model head enhances numerical precision by leveraging input from the autoregressive module, which simplifies the diffusion process by focusing on the next token.
  • Flexible generation: The autoregressive framework enables the generation of sequences with varying lengths, accommodating diverse scientific tasks.
  • Multimodality: By combining diffusion models with autoregressive modeling, UniGenX supports flexible multimodal data generation, bridging sequence and structure information.
  • State-of-the-Art Performance: Preliminary results demonstrate that UniGenX achieves significant improvements over existing methods across multiple scientific benchmarks.
Existing models often jointly diffuse sequence and structure or use sequences as embeddings. These approaches exacerbate training difficulties or fail to model sequences adequately. UniGenX handles long sequences effectively, demonstrating superior performance in crystal structure prediction for large systems.
We evaluate UniGenX on materials and small molecules, spanning tasks like crystal structure prediction, de novo material generation, conformation generation, and conditional molecule generation. UniGenX significantly outperforms FlowMM [39] on MP-20, Carbon-24, and MPTS-52 benchmarks and achieves state-of-the-art results in small molecule tasks. Trained on a unified dataset, UniGenX demonstrates cross-domain generalization. Incorporating pretraining with NatureLM [35] extends to natural language processing, validating our approach’s versatility.
In summary, UniGenX is a unified generative framework for science, integrating sequence and structure generation. By combining autoregressive and diffusion models, it addresses key limitations, enabling accurate and flexible scientific data generation.
2 Background
2.1 Autoregressive Model
$\mathbf{x}={x_{1},x_{2},\ldots,x_{T}},$
the auto-regressive approach models the probability as
$P(\mathbf{x})=\prod_{t=1}^{T}P(x_{t}\mid x_{1},x_{2},\ldots,x_{t-1}),$
At the heart of these transformer [28] decoder architectures lies the multi-head attention mechanism. Multi-head attention works by projecting an input sequence into multiple subspaces, allowing the model to attend to different parts of the sequence simultaneously. For a given token, the mechanism computes attention weights that quantify how much focus should be given to each other token in the sequence. These weights are then used to aggregate information from the sequence, creating a context-aware representation of the token.
Specifically, for an input sequence represented as a matrix $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$ (where $T$ is the sequence length and $d_{\text{model}}$ is the model’s dimensionality), the input embeddings are projected into $h$ independent subspaces, where $h$ is the number of attention heads. This is achieved by using separate learned weight matrices for each $i$-th head:
$\mathbf{Q}_{i}=\mathbf{X}\mathbf{W_{Q}}^{(i)},\quad\mathbf{K}_{i}=\mathbf{X}\mathbf{W_{K}}^{(i)},\quad\mathbf{V}_{i}=\mathbf{X}\mathbf{W_{V}}^{(i)},\quad i=1,2,\ldots,h.$
Here, $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$, $\mathbf{V}_{i}$ are called query, key and value respectively. $\mathbf{W_{Q}}^{(i)}$, $\mathbf{W_{K}}^{(i)}$, and $\mathbf{W_{V}}^{(i)}\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$ are the learned weight matrices for the $i$-th attention head, and $d_{k}$ is typically set to $d_{\text{model}}/h$ so that the total computational cost remains constant.
The attention weights are calculated as scaled dot-products between the queries and keys, followed by a softmax operation to normalize them:
$\text{Attention}(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i})=\text{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}_{i},$
where Attention $i\in\mathbb{R}^{T\times d_{k}}$ is the output of the $i$-th attention head, and the division by $\sqrt{d_{k}}$ serves as a scaling factor to prevent the dot product values from becoming too large, which could otherwise lead to small gradients during backpropagation.
The final output is calculated by concatenating the outputs for $h$ heads along the feature dimension and followed by a linear transformation:
$\mathbf{O}=\text{Concat}(\text{Attention}_{1},\text{Attention}_{2},\ldots,\text{Attention}_{h})\cdot\mathbf{W_{O}},$
where $\text{Concat}(\cdot)$ denotes the concatenation operation, $\mathbf{W_{O}}\in\mathbb{R}^{(h\cdot d_{k})\times d_{\text{model}}}$. The resulting matrix $\mathbf{O}\in\mathbb{R}^{T\times d_{\text{model}}}$ represents the final output of the multi-head attention mechanism, which is then passed to subsequent layers in the transformer architecture.
The transformer decoder architecture applies multi-head attention in an autoregressive manner by masking future tokens during training. This is achieved by using a causal mask, which prevents the model from attending to tokens that come after the current token in the sequence. This masking ensures that predictions are conditioned only on previously observed tokens, preserving the autoregressive property.
2.2 Diffusion Models
2.2.1 Diffusion Process
$\mathrm{d}x_{t}=f(x_{t},t)\mathrm{d}t+g(t)\mathrm{d}\bm{B}_{t},$ (1)
where $\bm{B}_{s}$ is a standard Wiener process, $f(\cdot,t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift coefficient of $x(t)$, and $g(\cdot):\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient of $x(t)$.
[…]
$x_{t}=\alpha_{t}x+\sigma_{t}\epsilon.$ (4)
Here $\alpha_{t}$ and $\sigma_{t}$ are linear and variance variable, depending on the time $t$; $\epsilon$ is a random variable. In the diffusion model, $\alpha_{t}$ and $\sigma_{t}$ are given by designed schemes for VE and VP cases [22]; the random variable $\epsilon$ is a standard Gaussian noise variable, i.e., $\epsilon\sim N(0,I)$. In our work, we use the DDPM setting for training and the VP scheme for sampling [21, 44]. We also test the diffusion choice of EDM [45]. We obtained similar performance leveraging DDPM and EDM respectively, and listed the results as an ablation study in Appendix E.3.2.
2.2.2 Diffusion Targets
$p(x_{t}|x)=Ce^{-\frac{-\|x_{t}-\alpha_{t}x\|^{2}}{2\sigma_{t}^{2}}}.$ (5)
[…]
where we finally obtain
[…]
$s_{\theta^{*}}(x_{t})$ $=-\frac{\epsilon_{\theta^{*}}(x_{t})}{\sigma_{t}}.$
2.3 Related Work
Recent research has explored the integration of autoregressive models with diffusion-based generative techniques to overcome the limitations of each approach and enhance overall generation capabilities. By leveraging diffusion models as a complementary mechanism within autoregressive frameworks, researchers aim to improve the precision and efficiency of tasks like continuous embedding generation, which traditionally posed challenges for autoregressive methods due to their reliance on vector quantization. One prominent line of research involves adding diffusion models as lightweight components on top of existing autoregressive language models. These models enhance precision in the generation of continuous representations by combining the strengths of both approaches. For example, the Masked Autoregressive (MAR) model [47] and LatentLLM [48] incorporate diffusion-based approaches to refine hidden representations and improve the generation of continuous embeddings. The NOVA model [49] exemplifies the success of this hybrid approach by integrating diffusion methods into an autoregressive framework for multimodal tasks. The result is a system that can handle complex generative tasks with higher precision and versatility, showcasing the potential of combining autoregressive and diffusion techniques to push the boundaries of generative modeling. These works are most closely related to ours, but require advanced encoders for other modalities (such as images) to generate hidden embeddings, followed by alignment with language in a latent space.
@@ 3.1 Sequentialize All for Multi-Domain/Task Compatibility
Figure 1: Sequentialization approach combines symbolic (words) and numerical data (numbers), separated by special tokens that identify the category of each data element. As illustrated in the figure for small molecules and materials, SMILES strings or chemical formulas are treated as words and delimited by special tokens such as $<$bos$>$ (beginning of SMILES) or $<$bof$>$ (beginning of formula). Corresponding coordinates are treated as numbers, delimited by tokens such as $<$boc$>$ (beginning of coordinates). This method provides a general framework applicable to diverse scientific data formats.
@@
To fully leverage the flexibility of multimodality, we propose a simple yet effective sequentialization approach: representing all data—formulas, coordinates, energies, forces, and so on—as a single sequence segmented by special tokens. This sequentialization enables framing all generation tasks as next-token prediction problems, thereby harnessing the scalability and flexibility inherent in autoregressive models for scientific applications. And we will show that this approach can handle different scientific domains with different tasks.
Figure 1 depicts our sequentialization scheme, which unifies symbolic (words) and numerical data (numbers) into a single sequence delineated by special tokens. For small molecules, the $<$bos$>$ token marks the onset of the SMILES string, while $<$boc$>$ signals the commencement of its coordinate sequence. A parallel approach is applied to materials, with $<$bof$>$ initiating the chemical formula and $<$boc$>$ the coordinate sequence. The $<$eos$>$ token is appended to signify the sequence’s termination. This methodology facilitates the encoding of diverse scientific data into a unified sequence format and is adaptable to other data types, including protein and DNA sequences.
This sequentialization method offers a simple and effective way to represent diverse scientific data across various domains and tasks. However, traditional sequence models struggle to handle numerical data (numbers) with the same efficacy as symbolic data (words), particularly when numerical precision is crucial. To address this challenge, our model incorporates a diffusion strategy combined with a sequence model that effectively handles both numerical and symbolic data within a sequence modeling framework. This approach will be detailed in the following subsection.
@@ 4 Experiments
To evaluate UniGenX, we conduct experiments across various tasks in the materials and small molecule domains. In this work, we present two model configurations of different sizes: UniGenX(100M) and UniGenX(400M), where the transformer decoder backbones contain 100M and 400M parameters, respectively. See Table S5 for detailed model configurations.

### v2

# UniGenX: a unified generative foundation model that couples sequence, structure and function to accelerate scientific design across proteins, molecules and materials

@@ Abstract
Function in natural systems arises from one-dimensional sequences forming three-dimensional structures with specific properties. However, current generative models suffer from critical limitations: training objectives seldom target function directly, discrete sequences and continuous coordinates are optimized in isolation, and conformational ensembles are under-modeled. We present UniGenX, a unified generative foundation model that addresses these gaps by co-generating sequences and coordinates under direct functional and property objectives across proteins, molecules, and materials. UniGenX represents heterogeneous inputs as a mixed stream of symbolic and numeric tokens, where a decoder-only autoregressive transformer provides global context and a conditional diffusion head generates numeric fields steered by task-specific tokens. Besides the new high SOTAs on structure prediction tasks, the model demonstrates state-of-the-art or competitive performance for the function-aware generation across domains: in materials, it achieves “conflicted” multi-property conditional generation, yielding 436 crystal candidates meeting triple constraints, including 11 with novel compositions and 4 confirmed thermodynamically stable by DFT; in chemistry, it sets new benchmarks on five property targets and conformer ensemble generation on GEOM; and in biology, it improves success in modeling protein induced fit (RMSD << 2 Å) by over 23-fold and enhances EC-conditioned enzyme design. Ablation studies and cross-domain transfer substantiate the benefits of joint discrete–continuous training, establishing UniGenX as a significant advance from prediction to controllable, function-aware generation.
Keywords— Unified Generation, Autoregressive Diffusion Model, Next-token Prediction, AI for Science
@@ 1 Introduction
The ability of one-dimensional sequences to form three-dimensional structures with specific functions is a fundamental principle in natural systems. Replicating this generative process is a central goal for artificial intelligence in scientific discovery, promising to unlock transformative applications in materials design, drug discovery, and protein engineering [1, 2]. The intricate interplay between a system’s discrete sequence (e.g., chemical formula, amino acid chain) and its continuous coordinates (e.g., atomic positions, protein fold) dictates its essential properties and functions. For instance, a crystal’s atomic arrangement governs its electronic characteristics [3], a molecule’s spatial conformation determines its pharmacological efficacy [4], and a protein’s 3D shape enables its biological role [5]. Given this sensitivity, where even minute deviations in atomic coordinates can nullify function, the challenge is not just to predict structures, but to generate them with high numerical precision under desired functional constraints.
However, current generative models face critical limitations that hinder this goal. First, most training objectives target structural accuracy in isolation, rather than the ultimate function or property of interest. Second, the generation of discrete sequences and continuous coordinates are often treated as separate tasks, failing to capture their co-dependent relationship. Finally, models typically generate a single static structure, under-modeling the dynamic conformational ensembles that are crucial for the function of many molecular and biological systems. These gaps have created a significant bottleneck, preventing a unified approach to controllable, function-aware generation across scientific domains.
The current modeling landscape is dominated by two powerful yet incomplete paradigms. On one hand, autoregressive language models (LLMs) excel at processing and generating discrete token sequences, demonstrating remarkable flexibility, scalability, and long-context reasoning [6, 7]. However, their token-based nature struggles to enforce the strict numerical precision required for generating physically plausible 3D atomic coordinates. On the other hand, diffusion models have achieved state-of-the-art performance in generating high-fidelity continuous data, such as images and 3D structures [8, 9, 10]. Yet, they are less adept at handling discrete sequence information and lack the flexible, token-by-token generative framework of autoregressive models, making the joint generation of sequence and structure challenging.
To address these shortcomings, we introduce UniGenX, a unified generative foundation model that co-generates sequences and coordinates under direct functional and property objectives. UniGenX bridges the gap between discrete and continuous generation by integrating an autoregressive transformer with a conditional diffusion head. In our framework, heterogeneous inputs are represented as a mixed stream of symbolic and numeric tokens. A decoder-only transformer processes this stream to capture global context, while the conditional diffusion head generates precise numeric fields (i.e., coordinates) steered by task-specific functional tokens. This hybrid architecture synergistically combines the contextual power and flexibility of autoregressive models with the numerical precision of diffusion models, enabling a single framework to operate across proteins, molecules, and materials.
UniGenX marks a significant advance from prediction to controllable, function-aware generation. Beyond setting new state-of-the-art benchmarks on structure prediction tasks, our model demonstrates unparalleled capabilities in conditional generation across domains. In materials science, it successfully navigates “conflicted” multi-property objectives to design 436 novel crystal candidates, including 11 with new compositions and 4 confirmed as thermodynamically stable by DFT calculations. In chemistry, it establishes new performance records on five property-targeting tasks and conformer ensemble generation. For biology, it improves the success rate of modeling protein induced fit (RMSD << 2 Å) by over 23-fold and enhances enzyme design conditioned on Enzyme Commission (EC) numbers. Through extensive ablation studies and cross-domain transfer experiments, we substantiate the benefits of our joint discrete–continuous training strategy, establishing UniGenX as a versatile and powerful foundation for AI-driven scientific discovery.
@@ 2.1 Sequentialize All for Multi-Domain/Task Compatibility
Figure 1: Sequentialization approach combines symbolic (words) and numerical data (numbers), separated by special tokens that identify the category of each data element. As illustrated in the figure for small molecules and materials, SMILES strings or chemical formulas are treated as words and delimited by special tokens such as <<bos>> (beginning of SMILES) or <<bof>> (beginning of formula). Corresponding coordinates are treated as numbers, delimited by tokens such as <<boc>> (beginning of coordinates). This method provides a general framework applicable to diverse scientific data formats.
@@
To fully leverage the flexibility of multimodality, we propose a simple yet effective sequentialization approach: representing all data—formulas, coordinates, energies, forces, and so on—as a single sequence segmented by special tokens. This sequentialization enables framing all generation tasks as next-token prediction problems, thereby harnessing the scalability and flexibility inherent in autoregressive models for scientific applications. And we will show that this approach can handle various tasks across different scientific domains.
Figure 1 depicts our sequentialization scheme, which unifies symbolic (words) and numerical data (numbers) into a single sequence delineated by special tokens. For small molecules, the <<bos>> token marks the onset of the SMILES string, while <<boc>> signals the commencement of its coordinate sequence. A parallel approach is applied to materials, with <<bof>> initiating the chemical formula and <<boc>> the coordinate sequence. The <<eos>> token is appended to signify the sequence’s termination. This methodology facilitates the encoding of diverse scientific data into a unified sequence format and is adaptable to other data types, including protein and DNA sequences. The explanation of the special tokens is provided in the Table S9.
This sequentialization method offers a simple and effective way to represent diverse scientific data across various domains and tasks. However, traditional sequence models struggle to handle numerical data (numbers) with the same efficacy as symbolic data (words), particularly when numerical precision is crucial. A diffusion strategy combined with a sequence model effectively handles both numerical and symbolic data within a sequence modeling framework. This approach will be detailed in the following subsection.
@@ 2.2 Bridge the Gap: Unified Sequence Modeling of Words and Numbers
Remark 1.
Benefits of Unified training of Sequence and Structure. Beyond the advantages for generation capabilities derived from the unified generation of sequence and structure, we emphasize the training perspective. Unified training allows the model to more comprehensively capture the inherent knowledge within scientific data. To illustrate, let $s$ denote the sequence, $x$ the coordinates of the structure, and $f_{\theta}$ the model with training parameters $\theta$. The learned mappings for different models are as follows:
$\begin{cases}f_{\theta}^{\text{diffusion}}:s\mapsto x,&\text{Sequence-to-structure diffusion models,}\\ f_{\theta}^{\text{sequence}}:s\mapsto s,&\text{Sequence-to-sequence autoregressive models,}\\ f_{\theta}^{\text{UniGenX{}}}:(s,x)\mapsto(s,x),&\text{UniGenX{} model.}\end{cases}$
The joint mapping $f_{\theta}^{\text{UniGenX{}}}$ from $(s,x)$ to $(s,x)$ significantly enhances knowledge distillation from scientific data compared to the other two mappings. In scientific domains, the properties or functions of a molecule are determined by both its sequence and structural coordinates, which are richly interconnected. The joint training strategy enables UniGenX to uncover these interdependencies, leading to improved performance in understanding and generation tasks. Consider the co-evolution of amino acids in a protein for example: two amino acids, though not adjacent in the sequence, may be spatially close in the 3D structure. Consequently, a mutation in one amino acid can induce a related mutation in the other. A sequence-only model would entirely miss this information, and a sequence-to-structure model, while capturing some information from extensive sequence-structure data, lacks explicit sequence modeling due to the absence of a sequence-based loss.
@@ 3 Experiments
To evaluate UniGenX, we conduct experiments across various tasks in the materials and small molecule domains. In this work, we present two model configurations of different sizes: UniGenX(100M) and UniGenX(400M), where the transformer decoder backbones contain 100M and 400M parameters, respectively. See Table S5 for detailed model configurations. In general, UniGenX(400M) outperfroms UniGenX(100M) reflecting scaling law in this model architecture.